OpenStudy (anonymous):

solve each system of equations: 2) x+y+z=0 2x+3y+2z= -1 x-y+z= 2

jhonyy9 (jhonyy9):

- assume first with 3rd x+y+z=0 x-y+z=2 --------- 2x+2z=2 so x+y=1 so x=1-y - multiply first by -2 and assume with second -2x-2y-2z=0 2x+3y+2z=-1 --------------- 0 y 0 = -1 so y=-1 than x=1-(-1)=1+1=2 so x=2 - than x+y+z=0 so 2-1+z=0 so 1+z=0 so z=-1 x=2 y=-1 z=-1